OpenStudy (anonymous):

Geometry: Finding endpoint. R(6,-2) and M(5,3)

OpenStudy (anonymous):

M is the midpoint?

OpenStudy (anonymous):

Yes, M is the midpoint.

OpenStudy (anonymous):

6 -2 5 3 4 8

OpenStudy (anonymous):

The answer is supposed to be (4,8)

OpenStudy (anonymous):

you can use a formula if you like. but i hope the pattern is clear. 6 to 5 one left then one more to 4 -2 to 3 is 5 up then 5 more to 8

OpenStudy (anonymous):

set up equations 5=(6+x)/2 3=(-2+y)/2 10=6+x so x=4 6=-2+y so y=8 (4,8)

OpenStudy (anonymous):

yeah and you can check that it works. the average of 4 and 6 is 5 and the average of -2 and 8 is 3 so midpoint of (6,-2) and (4,8) is (5,3)